OpenStudy (anonymous):

Find the area bounded by the curves y1 = x , y2 = − x , and y3 = −x + 2

OpenStudy (anonymous):

none of those are curves hehe

OpenStudy (amistre64):

as far as i can tell; those dont bind to each other

OpenStudy (anonymous):

no such area. y2 = -x and y3 = -x + 2 are parallel

OpenStudy (anonymous):

ooops it is sqrt of x and sqrt of -x

OpenStudy (anonymous):

Is y3 still -x + 2?

OpenStudy (amistre64):

y=sqrt(x) y=- sqrt(x) y = -x+2 will create an area

OpenStudy (amistre64):

there is no such thing as sqrt(-x) in real numbers lol

OpenStudy (anonymous):

\[y1=\sqrt{x}, y2=-\sqrt{x}, y3=-x+2\]

OpenStudy (amistre64):

we can flip this to its inverse as well; y = x^2; y = -x+2 and integrate from there as well

OpenStudy (amistre64):

x^2 = -x+2 x^2 +x - 2 = 0 x = -2, 1

OpenStudy (anonymous):

of course there is \[\sqrt{-x}\] domain is non-positive numbers!

OpenStudy (amistre64):

the area of -x+2 from -2 to 0 = [S] [-x+2] - [x^2] dx ; [-2,0]

OpenStudy (amistre64):

sqrt(-x) is complex imaginary numbers....

OpenStudy (amistre64):

-x^2/2 +2x - x^3/3 ; [-2,1]

OpenStudy (anonymous):

wait... how did you get x^2 if you squared the \[\sqrt{x}\]

OpenStudy (anonymous):

do you know how to invert a function?\[f \rightarrow f^{-1}\]

OpenStudy (amistre64):

-1/2 + 2 - 1/3 is one part of it: 7/6 -4/2 - 4 + 8/3 = -10/3 -10/3 - 7/6 = -27/6 area = 27/6....if i did it right :)

OpenStudy (amistre64):

y = sqrt(x) and y = -sqrt(x) inverts to y = x^2 :)

OpenStudy (anonymous):

oh ok

OpenStudy (amistre64):

y = -x+2; ionverts to y = -x+2 lol

OpenStudy (anonymous):

i disagree. no one says -x is negative. In any cases if you draw the picture and then rotate 90 degrees clockwise you see \[y=x^2\] and \[y=x+2\] they intersect at -1 and 2 so integrate \[\int^2_{-1} x+2-x^2\,dx\]

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

i got that the 3 bound are 0, 1, 4

OpenStudy (anonymous):

You want the intersection of \[x^2\] and \[y=x+2\] \[x^2=x+2\] \[x^2-x-2=0\] \[(x-2)(x+1)=0\] \[x=-1, x=2\]

OpenStudy (anonymous):

but none of the curves touch at -1.

OpenStudy (amistre64):

the intersection of x^2 and -x+2 was the original

OpenStudy (amistre64):

its odd that you disagree that sqrt(-x) is not a real number.....

OpenStudy (amistre64):

lovhap; did you get a right answer :)